# Optimal Control of a Subdiffusion model using Dirichlet-Neumann and Neumann-Neumann Waveform Relaxation Algorithms

### v1

@@ 7 Convergence of NNWR Algorithm
(40) $\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}p_{i}^{(k)}+y_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ p_{i}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;(x_{i-1},x_{i}),\\ p_{i}^{(k)}=\upsilon_{i-1}^{(k-1)},&\textrm{at }x_{i-1},i\neq 1,\\ p_{i}^{(k)}=\upsilon_{i}^{(k-1)},&\textrm{at }x_{i},i\neq N,\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}\chi_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}\chi_{i}^{(k)}+\psi_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ \chi_{i}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;(x_{i-1},x_{i}),\\ -\kappa_{i}\frac{d}{dx}\chi_{i}^{(k)}=\kappa_{i-1}\frac{d}{dx}p_{i-1}^{(k)}+\kappa_{i}\frac{d}{dx}p_{i}^{(k)},&\textrm{at }x_{i-1},i\neq 1,\\ \kappa_{i}\frac{d}{dx}\chi_{i}^{(k)}=\kappa_{i}\frac{d}{dx}p_{i}^{(k)}+\kappa_{i+1}\frac{d}{dx}p_{i+1}^{(k)},&\textrm{at }x_{i},i\neq N.\end{cases}\end{array}$
The formula is then used to update the interface values for $i=1,\ldots,N-1$ becomes:
@@
After discretization the system becomes:
(43) $\begin{array}[]{rcll}\begin{cases}AX_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}X_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ X_{i}^{(k)}=\Pi_{i-1}^{(k-1)},&\textrm{at }x_{i-1},i\neq 1,\\ X_{i}^{(k)}=\Pi_{i}^{(k-1)},&\textrm{at }x_{i},i\neq N,\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}A\Xi_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}\Xi_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ -\kappa_{i}\frac{d}{dx}\Xi_{i}^{(k)}=\kappa_{i-1}\frac{d}{dx}X_{i-1}^{(k)}+\kappa_{i}\frac{d}{dx}X_{i}^{(k)},&\textrm{at }x_{i-1},i\neq 1,\\ \kappa_{i}\frac{d}{dx}\Xi_{i}^{(k)}=\kappa_{i}\frac{d}{dx}X_{i}^{(k)}+\kappa_{i+1}\frac{d}{dx}X_{i+1}^{(k)},&\textrm{at }x_{i},i\neq N.\end{cases}\end{array}$
The formula is then used to update the interface values for $i=1,\ldots,N-1$ becomes:
@@
A quick calculation of the Dirichlet subproblems in (43) for $i=1,\ldots,N$ gives:
@@
And using the solutions (45), Neumann subproblems in (43) give:
$\hat{\Xi}^{(k)}_{1}(x,s)$ $=P\sinh\left((x-x_{0})\sqrt{\Lambda/\kappa_{1}}\right)C^{(k-1)}_{1},$
$\hat{\Xi}^{(k)}_{i}(x,s)$ $=P\cosh\left((x-x_{i-1})\sqrt{\Lambda/\kappa_{i}}\right)C^{(k-1)}_{i}+P\cosh\left((x_{i}-x)\sqrt{\Lambda/\kappa_{i}}\right)D^{(k-1)}_{i},\mbox{for $i=2,\ldots,N-1,$}$
$\hat{\Xi}^{(k)}_{N}(x,s)$ $=P\sinh\left((x_{N}-x)\sqrt{\Lambda/\kappa_{N}}\right)D^{(k-1)}_{N}.$
To make things simpler, we define $\sigma_{i}:=\sinh\left(a_{i}\sqrt{\Lambda/\kappa_{i}}\right)$ and $\gamma_{i}:=\cosh\left(a_{i}\sqrt{\Lambda/\kappa_{i}}\right)$. Then putting the boundary values in Neumann subproblems, we have:
@@
$C^{(k-1)}_{i}$ $=\sigma_{i}^{-1}\left(\left(\gamma_{i}\sigma_{i}^{-1}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\gamma_{i+1}\sigma_{i+1}^{-1}\right)P^{-1}\Pi_{i}^{(k-1)}-\sigma_{i}^{-1}P^{-1}\Pi_{i-1}^{(k-1)}-\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\sigma_{i+1}^{-1}P^{-1}\Pi_{i+1}^{(k-1)}\right),\quad i=2,\ldots,N-1,$
$D^{(k-1)}_{i}$ $=\sigma_{i}^{-1}\left(\left(\gamma_{i}\sigma_{i}^{-1}+\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\gamma_{i-1}\sigma_{i-1}^{-1}\right)P^{-1}\Pi_{i-1}^{(k-1)}-\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\sigma_{i-1}^{-1}P^{-1}\Pi_{i-2}^{(k-1)}-\sigma_{i}^{-1}P^{-1}\Pi_{i}^{(k-1)}\right),\quad i=2,\ldots,N-1,$
@@
Substituting $\Xi_{i}$ values in updating equation (44) and using the identity $\gamma_{i}^{2}-\sigma_{i}^{2}=I$, we have
@@
Using $\theta_{i}=\bar{\theta}_{i}:=1/(2+\sqrt{\kappa_{i}/\kappa_{i+1}}+\sqrt{\kappa_{i+1}/\kappa_{i}})$ and $\phi_{i}=\bar{\phi}_{i}:=1/(2+\sqrt{\kappa_{i}/\kappa_{i+1}}+\sqrt{\kappa_{i+1}/\kappa_{i}})$for super-linear convergence in (46)-(48), we have:
@@
Lemma 7.1.
@@
Proof 7.2.
The procedure to prove all five parts is the same, with the only difference being the combination of exponentials. Thus, we will present the proof up to certain standard levels for each case, after which they will follow the same procedure.
@@
    Rewrite the hyperbolic terms in exponential we obtain:
    $\left|\frac{\cosh(l_{1}z)\cosh(l_{2}z)}{\sinh(l_{1}z)\sinh(l_{2}z)}-1\right|$
@@
    $\leq 2\left|\frac{\exp(-2l_{1}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|+2\left|\frac{\exp(-2l_{2}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|$
    Now for the first part, we get by part (i) of the Lemma LABEL:SimpleLaplaceLemma
    $\left|\frac{\exp(-2l_{1}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|$ $\leq\left|\frac{\exp(-l_{1}z)}{(1-\exp(-2l_{1}z))}\right|\left|\frac{\exp(-l_{1}z)}{(1-\exp(-2l_{2}z))}\right|$
@@
    $=\frac{\exp(-2l_{1}x_{0})}{(1-\exp(-2l_{1}x_{0}))(1-\exp(-2l_{2}x_{0}))}$
    similarly,
    $\left|\frac{\exp(-2l_{2}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|\leq\frac{\exp(-2l_{2}x_{0})}{(1-\exp(-2l_{1}x_{0}))(1-\exp(-2l_{2}x_{0}))}$
    Therefore
@@
    Similar simplification is possible for the second part $I_{2}$. Finally, we have our result using Lemma LABEL:estimate_1 .
    Proofs of parts (ii) and (iii) are similar to part (i), hence omitted.
@@
    Using the triangular inequality, we have:
    $\left|\frac{1}{\sinh(l_{0}z)}\left(\rho_{1}\frac{\cosh(l_{1}z)}{\sinh(l_{1}z)}-\rho_{2}\frac{\cosh(l_{2}z)}{\sinh(l_{2}z)}\right)\right|\leq\rho_{1}\left|\frac{\cosh(l_{1}z)}{\sinh(l_{0}z)\sinh(l_{1}z)}\right|+\rho_{2}\left|\frac{\cosh(l_{2}z)}{\sinh(l_{0}z)\sinh(l_{2}z)}\right|$
    Now rewrite the hyperbolic terms in exponential for $I_{3}$ give:
@@
    $\leq\rho_{1}\frac{\cosh(l_{1}x_{0})}{\sinh(l_{0}x_{0})\sinh(l_{1}x_{0})}$
    Similarly
    $\rho_{2}\left|\frac{\cosh(l_{2}z)}{\sinh(l_{0}z)\sinh(l_{2}z)}\right|\leq\rho_{2}\frac{\cosh(l_{2}x_{0})}{\sinh(l_{0}x_{0})\sinh(l_{2}x_{0})}$
    Therefore
    $\left|\frac{1}{\sinh(l_{0}z)}\left(\rho_{1}\frac{\cosh(l_{1}z)}{\sinh(l_{1}z)}-\rho_{2}\frac{\cosh(l_{2}z)}{\sinh(l_{2}z)}\right)\right|\leq\frac{1}{\sinh(l_{0}x_{0})}\left(\rho_{1}\frac{\cosh(l_{1}x_{0})}{\sinh(l_{1}x_{0})}+\rho_{2}\frac{\cosh(l_{2}x_{0})}{\sinh(l_{2}x_{0})}\right)$
    similar for $I_{4}$. Finally, using Lemma LABEL:estimate_1 gives the result.
@@
    The proof is similar to part (iv) and thus omitted.
Now we are prepared to present the following convergence theorems:
Theorem 7.3 (Convergence of NNWR for sub-diffusion and diffusion-wave).
For $\theta_{i}=\bar{\theta}_{i}=1/(2+\sqrt{\kappa_{i}/\kappa_{i+1}}+\sqrt{\kappa_{i+1}/\kappa_{i}})$, the NNWR algorithm for the fractional order $(0<\nu\leq 1)$ in (LABEL:NNWRD)-(11) for time window $T$, follow the following bound:
$\max_{1\leq i\leq N-1}||\Pi_{i}^{(k)}||_{\infty}\leq\bar{d}^{k}\operatorname{cosech}^{2k}\left(\frac{h}{2}\lambda\right)\|P\|_{\infty}\|P^{-1}\|_{\infty}\max_{1\leq j\leq N-1}||\Pi_{i}^{(0)}||_{\infty}.$
Proof 7.4.
Define $h_{i}:=a_{i}/\sqrt{\kappa_{i}}$. Set $\hat{\upsilon}^{(k)}_{i}(s):=\sigma^{k}\hat{\omega}^{(k)}_{i}(s)$ to reduce the equations (LABEL:NN4)-(LABEL:NN6) to the form
@@
$T_{i,i}=\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\gamma_{i}\sigma_{i}^{-1}\gamma_{i+1}\sigma_{i+1}^{-1}-I\right),\,T_{i,i+1}=\sigma_{i+1}^{-1}\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\gamma_{i+2}\sigma_{i+2}^{-1}-\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\gamma_{i}\sigma_{i}^{-1}\right),$
@@
$T_{N-1,N-1}=\left(\sqrt{\frac{\kappa_{N}}{\kappa_{N-1}}}(\gamma_{N-1}\sigma_{N-1}^{-1}\gamma_{N}\sigma_{N}^{-1}-I)+\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}(\sigma_{N}\gamma_{N}^{-1}\gamma_{N-1}\sigma_{N-1}^{-1}-I)\right),$
@@
These weights $\hat{t}_{i,j}$ in (LABEL:NNWR3) are in the same form as in Lemma LABEL:lam_7. Therefore, we have the following estimates:
for $i=1,2,\cdots,N-1$,
@@
After having the inverse Laplace transform in (LABEL:NN7) and using part (ii) of Lemma LABEL:SimpleLaplaceLemma we obtain
$||P^{-1}\Pi_{1}^{(k)}||_{\infty}$ $\leq c_{1}\bar{\theta}_{1}\max_{1\leq j\leq N-1}||P^{-1}\Pi_{j}^{(k-1)}||_{\infty},$
(54) $||P^{-1}\Pi_{i}^{(k)}||_{\infty}$ $\leq c_{i}\bar{\theta}_{i}\max_{1\leq j\leq N-1}||P^{-1}\Pi_{j}^{(k-1)}||_{\infty},$
$||P^{-1}\Pi_{N-1}^{(k)}||_{\infty}$ $\leq c_{N-1}\bar{\theta}_{N-1}\max_{1\leq j\leq N-1}||P^{-1}\Pi_{j}^{(k-1)}||_{\infty}.$
@@
Define
(55) $\bar{c}:=\max_{1\leq i\leq N-1}\bar{\theta}_{i}c_{i}.$
Then from equations (LABEL:NNWR4), we have
(56) $||P^{-1}\Pi_{i}^{(k)}||_{\infty}\leq\bar{c}\max_{1\leq j\leq N-1}||P^{-1}\Pi_{i}^{(k-1)}||_{\infty},$
which further yields
$\max_{1\leq i\leq N-1}||\Pi_{i}^{(k)}||_{\infty}\leq\bar{c}^{k}\|P\|_{\infty}\|P^{-1}\|_{\infty}\max_{1\leq j\leq N-1}||\Pi_{i}^{(0)}||_{\infty},$
we can find the bound in terms of minimum subdomain length. For that:
$w_{i,i}$ $\leq\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\coth(h_{i}\lambda)\coth(h_{i+1}\lambda)-1\right)$
@@
$=\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\operatorname{cosech}^{2}(h\lambda)$
@@
$w_{i,i+1}$ $\leq\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\operatorname{cosech}(h\lambda)\coth(h\lambda)$
$w_{i,i+2}$ $\leq\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\operatorname{cosech}^{2}(h\lambda)$
$w_{i,i-1}$ $\leq\left(\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}+\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}\right)\operatorname{cosech}(h\lambda)\cot(h\lambda)$
$w_{i,i-2}$ $\leq\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\operatorname{cosech}^{2}(h\lambda).$
Therefore
$c_{1}$ $\leq\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}+\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\right)\operatorname{cosech}(h\lambda)(\operatorname{cosech}(h\lambda)+\coth(h\lambda))$
@@
$=\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}+\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\right)\frac{\operatorname{cosech}^{2}(\frac{h}{2}\lambda)}{2}$
$=:d_{1}\operatorname{cosech}^{2}\left(\frac{h}{2}\lambda\right)$
Similarly

### v2

@@ 8 Convergence of Neumann-Neumann Waveform Relaxation
(40) $\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}p_{i}^{(k)}+y_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ p_{i}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;(x_{i-1},x_{i}),\\ p_{i}^{(k)}=\upsilon_{i-1}^{(k-1)},&\textrm{at }x_{i-1},i\neq 1,\\ p_{i}^{(k)}=\upsilon_{i}^{(k-1)},&\textrm{at }x_{i},i\neq N,\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}\chi_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}\chi_{i}^{(k)}+\psi_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ \chi_{i}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;(x_{i-1},x_{i}),\\ -\kappa_{i}\frac{d}{dx}\chi_{i}^{(k)}=\kappa_{i-1}\frac{d}{dx}p_{i-1}^{(k)}+\kappa_{i}\frac{d}{dx}p_{i}^{(k)},&\textrm{at }x_{i-1},i\neq 1,\\ \kappa_{i}\frac{d}{dx}\chi_{i}^{(k)}=\kappa_{i}\frac{d}{dx}p_{i}^{(k)}+\kappa_{i+1}\frac{d}{dx}p_{i+1}^{(k)},&\textrm{at }x_{i},i\neq N,\end{cases}\end{array}$
with the update conditions at interfaces for $i=1,\ldots,N-1$:
@@
After employing the same discretization technique for the time fractional derivative, which encompasses both uniform and two-sided graded meshes as outlined in section 7, the NNWR system can be readily expressed as:
(43) $\begin{array}[]{rcll}\begin{cases}\mathbb{L}X_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}X_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ X_{i}^{(k)}=\Pi_{i-1}^{(k-1)},&\textrm{at }x_{i-1},i\neq 1,\\ X_{i}^{(k)}=\Pi_{i}^{(k-1)},&\textrm{at }x_{i},i\neq N,\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}\mathbb{L}\Xi_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}\Xi_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ -\kappa_{i}\frac{d}{dx}\Xi_{i}^{(k)}=\kappa_{i-1}\frac{d}{dx}X_{i-1}^{(k)}+\kappa_{i}\frac{d}{dx}X_{i}^{(k)},&\textrm{at }x_{i-1},i\neq 1,\\ \kappa_{i}\frac{d}{dx}\Xi_{i}^{(k)}=\kappa_{i}\frac{d}{dx}X_{i}^{(k)}+\kappa_{i+1}\frac{d}{dx}X_{i+1}^{(k)},&\textrm{at }x_{i},i\neq N.\end{cases}\end{array}$
The formula, subsequently employed to update the interface values for $i=1,2,\ldots,N-1$, be:
@@
After evaluating the solution of Dirichlet subproblems outlined in equation (43) for $i=1,\ldots,N$, we obtain:
@@
And using this Dirichlet solutions (45), Neumann subproblems in (43) give:
$\hat{\Xi}^{(k)}_{1}(x)$ $=P\sinh\left((x-x_{0})\sqrt{\Lambda/\kappa_{1}}\right)C^{(k-1)}_{1}$,
$\hat{\Xi}^{(k)}_{i}(x)$ $=P\cosh\left((x-x_{i-1})\sqrt{\Lambda/\kappa_{i}}\right)C^{(k-1)}_{i}+P\cosh\left((x_{i}-x)\sqrt{\Lambda/\kappa_{i}}\right)D^{(k-1)}_{i},\mbox{for $i=2,\ldots,N-1,$}$
$\hat{\Xi}^{(k)}_{N}(x)$ $=P\sinh\left((x_{N}-x)\sqrt{\Lambda/\kappa_{N}}\right)D^{(k-1)}_{N}$.
For the sake of simplifying notations, we consider: $\sigma_{i}:=\sinh\left(a_{i}\sqrt{\Lambda/\kappa_{i}}\right)$ and $\gamma_{i}:=\cosh\left(a_{i}\sqrt{\Lambda/\kappa_{i}}\right)$, and using boundary conditions in Neumann subproblems to evaluate the ode constants, which for $i=2,\ldots,N-1$ are:
@@
$C^{(k-1)}_{i}$ $=\sigma_{i}^{-1}\left(\left(\gamma_{i}\sigma_{i}^{-1}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\gamma_{i+1}\sigma_{i+1}^{-1}\right)P^{-1}\Pi_{i}^{(k-1)}-\sigma_{i}^{-1}P^{-1}\Pi_{i-1}^{(k-1)}-\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\sigma_{i+1}^{-1}P^{-1}\Pi_{i+1}^{(k-1)}\right),$
$D^{(k-1)}_{i}$ $=\sigma_{i}^{-1}\left(\left(\gamma_{i}\sigma_{i}^{-1}+\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\gamma_{i-1}\sigma_{i-1}^{-1}\right)P^{-1}\Pi_{i-1}^{(k-1)}-\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\sigma_{i-1}^{-1}P^{-1}\Pi_{i-2}^{(k-1)}-\sigma_{i}^{-1}P^{-1}\Pi_{i}^{(k-1)}\right),$
@@
By substituting the values of $\Xi_{i}$ into the updating part (44) and employing $\gamma_{i}^{2}-\sigma_{i}^{2}=I$, where $I$ imply identity, we obtain:
@@
Using $\theta_{i}=\phi_{i}=\bar{\theta}_{i}:=1/(2+\sqrt{\kappa_{i}/\kappa_{i+1}}+\sqrt{\kappa_{i+1}/\kappa_{i}})$ to eliminate the linear terms in (46)-(48), we have:
@@
Before going to the convergence result we need the subsequent lemma for estimation.
Lemma 14.
@@
Proof.
The procedure for proving all five parts is similar, with the only distinction being the combination of exponentials.
@@
    Rewriting the hyperbolic functions in terms of exponential than using the inequality yield
    (52) $\left|\frac{\cosh(l_{1}z)\cosh(l_{2}z)}{\sinh(l_{1}z)\sinh(l_{2}z)}-1\right|$
@@
    $\leq 2\left|\frac{\exp(-2l_{1}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|+2\left|\frac{\exp(-2l_{2}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|.$
    Now for the first part, using the Lemma 12 we have
    (53) $\left|\frac{\exp(-2l_{1}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|$ $\leq\left|\frac{\exp(-l_{1}z)}{(1-\exp(-2l_{1}z))}\right|\left|\frac{\exp(-l_{1}z)}{(1-\exp(-2l_{2}z))}\right|$
@@
    $=\frac{\exp(-2l_{1}x_{0})}{(1-\exp(-2l_{1}x_{0}))(1-\exp(-2l_{2}x_{0}))}.$
    In similar manner from the second part we obtain the estimate
    (54) $\left|\frac{\exp(-2l_{2}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|\leq\frac{\exp(-2l_{2}x_{0})}{(1-\exp(-2l_{1}x_{0}))(1-\exp(-2l_{2}x_{0}))}.$
    Therefore combining (53) and (54) in (52) provides
@@
    Proofs for parts (ii) and (iii) follow a similar approach as part (i), and are therefore omitted.
@@
    By employing the triangle inequality gives
    $\left|\frac{1}{\sinh(l_{0}z)}\left(\rho_{1}\frac{\cosh(l_{1}z)}{\sinh(l_{1}z)}-\rho_{2}\frac{\cosh(l_{2}z)}{\sinh(l_{2}z)}\right)\right|\leq\rho_{1}\left|\frac{\cosh(l_{1}z)}{\sinh(l_{0}z)\sinh(l_{1}z)}\right|+\rho_{2}\left|\frac{\cosh(l_{2}z)}{\sinh(l_{0}z)\sinh(l_{2}z)}\right|.$
    Rewrite the hyperbolic expressions in exponential form, then applying Lemma 12 to the initial part yields
@@
    $\leq\rho_{1}\frac{\cosh(l_{1}x_{0})}{\sinh(l_{0}x_{0})\sinh(l_{1}x_{0})}.$
    In similar manner we obtain the second part’s estimate, which is
    $\rho_{2}\left|\frac{\cosh(l_{2}z)}{\sinh(l_{0}z)\sinh(l_{2}z)}\right|\leq\rho_{2}\frac{\cosh(l_{2}x_{0})}{\sinh(l_{0}x_{0})\sinh(l_{2}x_{0})}.$
    Therefore combining both parts we obtain
    $\left|\frac{1}{\sinh(l_{0}z)}\left(\rho_{1}\frac{\cosh(l_{1}z)}{\sinh(l_{1}z)}-\rho_{2}\frac{\cosh(l_{2}z)}{\sinh(l_{2}z)}\right)\right|\leq\frac{1}{\sinh(l_{0}x_{0})}\left(\rho_{1}\frac{\cosh(l_{1}x_{0})}{\sinh(l_{1}x_{0})}+\rho_{2}\frac{\cosh(l_{2}x_{0})}{\sinh(l_{2}x_{0})}\right).$
@@
    The proof follows a similar approach as part (iv) and is consequently omitted.
Based on the aforementioned estimates, we are now able to introduce the principal convergent outcome of the NNWR algorithm.
Theorem 15 (Convergence of NNWR).
Let time discretize combine state and adjoint matrix $\mathbb{L}$ in (43) be diagonalizable by the eigenvector matrix $P$ corresponding to the eigenvalues $\{\lambda_{i}:i=1,\ldots,2n+2\}$. If $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$, then the interface error term of the NNWR algorithm in (9)-(10) in 1D for relaxation parameters $\theta_{i}=\phi_{i}=\bar{\theta}_{i}=1/(2+\sqrt{\kappa_{i}/\kappa_{i+1}}+\sqrt{\kappa_{i+1}/\kappa_{i}})$ satisfies the following bound
[…]
where $h=\min_{1\leq i\leq N}a_{i}/\sqrt{\kappa_{i}}$ and $\bar{d}=\max_{1\leq i\leq N-1}\bar{\theta}_{i}d_{i}$, the values of $d_{i}$ are given in (63) and (64).
Proof.
First we use the same argument as in the proof of Theorem 13 to show that $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$ is well defined. Then left multiplying on both side with $P^{-1}$ of the update terms in (49)-(51) to obtain
@@
(57) $T_{i,i}=\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\gamma_{i}\sigma_{i}^{-1}\gamma_{i+1}\sigma_{i+1}^{-1}-I\right),\,T_{i,i+1}=\sigma_{i+1}^{-1}\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\gamma_{i+2}\sigma_{i+2}^{-1}-\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\gamma_{i}\sigma_{i}^{-1}\right),$
@@
(58) $T_{N-1,N-1}=\left(\sqrt{\frac{\kappa_{N}}{\kappa_{N-1}}}(\gamma_{N-1}\sigma_{N-1}^{-1}\gamma_{N}\sigma_{N}^{-1}-I)+\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}(\sigma_{N}\gamma_{N}^{-1}\gamma_{N-1}\sigma_{N-1}^{-1}-I)\right),$
@@
These weights $T_{i,j}$ in (56)-(58) are in the same form as in Lemma 14. Therefore, using the notation $h_{j}=\frac{a_{j}}{\sqrt{\kappa_{j}}}:j=1,\ldots,N$, we obtain the subsequent estimations for $i=1,2,\cdots,N-1$:
@@
Now taking the supremum norm on both side in (55) we have
[…]
@@
Now we can find the weight bounds in terms of minimum subdomain length defined as $h:=\min_{1\leq i\leq N}h_{i}$. Utilizing the monotonically decreasing nature of the $\coth$ and $\operatorname{cosech}$ functions with positive arguments. The estimates are as follows:
(61) $w_{i,i}$ $=\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\coth(h_{i}\lambda)\coth(h_{i+1}\lambda)-1\right)$
@@
$=\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\operatorname{cosech}^{2}(h\lambda),$
@@
(62) $w_{i,i+1}$ $\leq\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\operatorname{cosech}(h\lambda)\coth(h\lambda),\,w_{i,i+2}\leq\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\operatorname{cosech}^{2}(h\lambda),$
$w_{i,i-1}$ $\leq\left(\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}+\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}\right)\operatorname{cosech}(h\lambda)\cot(h\lambda),\,w_{i,i-2}\leq\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\operatorname{cosech}^{2}(h\lambda).$
Therefore using the estimates from (61)-(62) in (8) we have
(63) $c_{1}$ $\leq\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}+\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\right)\operatorname{cosech}(h\lambda)(\operatorname{cosech}(h\lambda)+\coth(h\lambda))$
@@
$=\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}+\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\right)\frac{\operatorname{cosech}^{2}(\frac{h}{2}\lambda)}{2}=:d_{1}\operatorname{cosech}^{2}\left(\frac{h}{2}\lambda\right),$
and similarly